# RM-R1: Reward Modeling as Reasoning

### v1

@@ Abstract
Reward modeling is essential for aligning large language models (LLMs) with human preferences, especially through reinforcement learning from human feedback (RLHF). To provide accurate reward signals, a reward model (RM) should stimulate deep thinking and conduct interpretable reasoning before assigning a score or a judgment. However, existing RMs either produce opaque scalar scores or directly generate the prediction of a preferred answer, making them struggle to integrate natural language critiques, thus lacking interpretability. Inspired by recent advances of long chain-of-thought (CoT) on reasoning-intensive tasks, we hypothesize and validate that integrating reasoning capabilities into reward modeling significantly enhances RM’s interpretability and performance. In this work, we introduce a new class of generative reward models – Reasoning Reward Models (ReasRMs) – which formulate reward modeling as a reasoning task. We propose a reasoning-oriented training pipeline and train a family of ReasRMs, RM-R1. The training consists of two key stages: (1) distillation of high-quality reasoning chains and (2) reinforcement learning with verifiable rewards. RM-R1 improves LLM rollouts by self-generating reasoning traces or chat-specific rubrics and evaluating candidate responses against them. Empirically, our models achieve state-of-the-art or near state-of-the-art performance of generative RMs across multiple comprehensive reward model benchmarks, outperforming much larger open-weight models (e.g., Llama3.1-405B) and proprietary ones (e.g., GPT-4o) by up to 13.8%. Beyond final performance, we perform thorough empirical analysis to understand the key ingredients of successful ReasRM training. To facilitate future research, we release six ReasRM models along with code and data at https://github.com/RM-R1-UIUC/RM-R1.
@@ 1 Introduction
Large language models (LLMs) are rapidly becoming more capable and thus playing increasingly important roles for a wide spectrum of applications that require understanding [12], generation [6] and complex decision-making [16, 22] capabilities. Reward models (RMs) play a critical role in LLM post-training, particularly in reinforcement learning with human feedback (RLHF) [25, 4], where they serve as scalable proxies for human evaluators. A growing body of research has focused on developing domain-specific reward models, such as mathematical verifiers [35, 10]. In this work, we are interested in advancing generalist reward modeling, where a reward model is expected to provide robust and accurate supervision across a wide range of domains. Generalist reward modeling poses greater challenges, as the evaluation criteria are more diverse and complex compared to specialized settings. However, developing strong generalist reward models is crucial for enabling broader, more versatile LLM applications.
Existing research on reward modeling can be broadly classified into two categories: (1) scalar-based reward models (Scalar RM)  [21] and (2) generative reward models (GenRM)  [40]. Scalar-based approaches frame reward modeling as a classification problem, typically training a sequence-level classifier on top of a language model. In contrast, generative-based approaches retain the original language model decoding head and leverage the model’s generative abilities to produce free-form pairwise judgments. While scalar-based methods are direct and often effective, they are opaque, offering no intermediate reasoning steps to justify the model’s decisions. This lack of transparency may limit their capacity to handle more challenging, reasoning-intensive preference tasks. On the other hand, although generative-based methods provide greater transparency by exposing intermediate reasoning traces, their reasoning is often superficial and unhelpful for reliable judgment, leading to suboptimal performance [24, 5].
@@
As illustrated in Figure 1, correct preference judgments require accurately perceiving the question, understanding the evaluation rubrics (i.e., the major aspects to evaluate) and constructing coherent, convincing arguments—closely mirroring how humans approach grading tasks. A coherent, contemplative, and logically grounded reasoning process could naturally support solving challenging pairwise preference tasks, particularly as reasoning domains already constitute a major component of modern reward modeling and we have witnessed the success of reasoning models in those domains [11]. Motivated by these observations, we explore the following central question:
@@
In this work, we unleash the reasoning potential of reward models and propose a new class of models: Reasoning Reward Models (ReasRMs). We find that integrating long reasoning chains during the judging process significantly enhances downstream reward model performance, consistently outperforming its non-reasoning counterparts (e.g., scalar RM and ordinary GenRM). We explore several strategies for adapting instruction-tuned language models into logically coherent ReasRMs. Notably, we observe that directly applying reinforcement learning with verifiable rewards (RLVR) to generalist reward modeling fails to fully unleash the model’s reasoning potentials for reward modeling. Through a series of studies, we design a reasoning training pipeline that introduces a stage of structured reasoning distillation prior to RLVR, ultimately resulting in the development of RM-R1. We observe that the distillation stage is critical, as RLVR alone often biases the model toward focusing on superficial features.
Starting from an instruction-tuned model, RM-R1 training leverages a novel Chain-of-Rubrics (CoR) prompting framework that elicits structured and coherent reasoning. Specifically, the model first classifies each input task into one of two categories: reasoning or chat. For chat tasks, the model generates a set of evaluation rubrics, justifications for the rubrics, and evaluations tailored to the specific question. For reasoning tasks, correctness is generally preferred, so we directly let the model first solve the problem itself before evaluating the pairwise model responses. This finer-grained task classification enables RLVR to fully exploit reasoning supervision for reasoning-heavy tasks while producing coherent, high-quality evaluation chains for chat tasks.
In addition, we explore how to directly adapt existing reasoning models into reward models. Since these models have already undergone substantial reasoning-focused distillation, we fine-tune them using RLVR without requiring additional distillation stages. Based on our training recipes, we release a family of ReasRM models ranging from 7B to 32B. Empirically, our reasoning reward models consistently yield highly interpretable and coherent reasoning traces, achieving better or comparative performance on RewardBench [18], RM-Bench [23], and RMB [43], outperforming 70B, 340B, GPT-4o, and Claude models by up to 13.8% in accuracy. Beyond final performance, we conduct extensive empirical analyses of RM-R1, including studies of its scaling effects, ablations of our training recipes, comparisons with non-reasoning baselines, training dynamics, and detailed case studies.
In summary, our main contributions are as follows:
@@
  • We design a training recipe based on reasoning-oriented distillation and RL that produces a set of reward models - RM-R1- that can outperform larger models by up to 13.8% in accuracy.
@@
  • We open-source the related resources, including code, data, and model checkpoints, to facilitate reproducibility and follow-up research.

### v2

@@ Abstract
Reward modeling is essential for aligning large language models (LLMs) with human preferences through reinforcement learning (RL). To provide accurate reward signals, a reward model (RM) should stimulate deep thinking and conduct interpretable reasoning before assigning a score or a judgment. Inspired by recent advances of long chain-of-thought (CoT) on reasoning-intensive tasks, we hypothesize and validate that integrating reasoning capabilities into reward modeling significantly enhances RM’s interpretability and performance. To this end, we introduce a new class of generative reward models – Reasoning Reward Models (ReasRMs) – which formulate reward modeling as a reasoning task. We propose a reasoning-oriented training pipeline and train a family of ReasRMs, RM-R1. RM-R1 features a chain-of-rubrics (CoR) mechanism – self-generating sample-level chat rubrics or math/code solutions, and evaluating candidate responses against them. The training of RM-R1 consists of two key stages: (1) distillation of high-quality reasoning chains and (2) reinforcement learning with verifiable rewards. Empirically, our models achieve state-of-the-art performance across three reward model benchmarks on average, outperforming much larger open-weight models (e.g., INF-ORM-Llama3.1-70B) and proprietary ones (e.g., GPT-4o) by up to 4.9%. Beyond final performance, we perform thorough empirical analysis to understand the key ingredients of successful ReasRM training. To facilitate future research, we release six ReasRM models along with code and data at https://github.com/RM-R1-UIUC/RM-R1.
@@ 1 Introduction
Reward models (RMs) play a critical role in LLM post-training, particularly in reinforcement learning with human feedback (RLHF) [4, 23], where they serve as scalable proxies for human evaluators. Existing research on reward modeling can be broadly classified into two categories: (1) scalar-based reward models (ScalarRM)  [19] and (2) generative reward models (GenRM)  [37]. Scalar-based approaches frame reward modeling as a classification problem, typically training a sequence classifier on top of a language model. In contrast, generative approaches retain the original language model decoding head and leverage the model’s generative abilities to produce free-form pairwise judgments. While scalar-based methods are direct and often effective, they are opaque, offering no intermediate reasoning steps to justify the model’s decisions. This lack of transparency may limit their capacity to handle more challenging, reasoning-intensive preference tasks. On the other hand, although generative methods provide greater transparency, their reasoning is often superficial and unhelpful for reliable judgment, leading to suboptimal performance [7, 21].
@@
In real-world decision-making scenarios, accurate and grounded reward modeling often requires jointly conducting reasoning and reward assignment. This is because preference judgments inherently involve multifaceted cognitive considerations, such as inferring a judge’s latent evaluation criteria [5], navigating trade-offs among multiple criteria [22], and simulating potential consequences [31], all of which necessitate extensive reasoning. Our example in Figure 1 illustrates such an example, where a correct preference judgement requires accurate perception of the question, understanding of the corresponding evaluation rubrics with convincing arguments – closely mirroring how humans approach grading tasks. Motivated by these observations, we explore the following central question:
@@
In this work, we unleash the reasoning potential of RMs and propose a new class of models: Reasoning Reward Models (ReasRMs). Different from standard GenRMs, ReasRMs emphasize leveraging long and coherent reasoning chains during the judging process to enhance the model’s ability to assess and distinguish complex outputs accurately. We hypothesize and validate that integrating long reasoning chains during the judging process significantly enhances downstream reward model performance. We explore several strategies for adapting instruction-tuned language models into logically coherent ReasRMs. Notably, we find that solely applying reinforcement learning with verifiable rewards (RLVR) [12] in reward modeling does not fully realize the model’s reasoning capabilities. We also observe that plain chain-of-thought (CoT) reasoning falls short at perceiving the fine-grained distinction across different question types.
Through a series of studies, we design a training pipeline that introduces reasoning distillation prior to RLVR, ultimately resulting in the development of RM-R1. To guide RM-R1 for structured rollouts tailored for reward modeling during RL training and inference, we design a Chain-of-Rubrics (CoR) process. Specifically, the model categorizes the input sample into one of two categories: chat or reasoning. For chat tasks, the model generates a set of evaluation rubrics, justifications for the rubrics, and evaluations tailored to the specific question. For reasoning tasks, correctness is the most important and generally preferred rubrics, so we directly let the model first solve the problem itself before evaluating and picking the preferred response. This task perception enables RLVR to fully exploit reasoning supervision for reasoning-heavy tasks while producing coherent, high-quality evaluation chains for chat tasks. In addition, we explore how to directly adapt existing reasoning models into reward models. Since these models have already undergone substantial reasoning-focused distillation, we fine-tune them using RLVR without additional distillation stages. Based on our training recipes, we produce RM-R1 models ranging from 7B to 32B.
Empirically, RM-R1 models consistently yield highly interpretable and coherent reasoning traces. On average, RM-R1 achieves state-of-the-art performance on RewardBench [16], RM-Bench [20], and RMB [39], outperforming 70B, 340B, GPT-4o, and Claude models by up to 4.9% in accuracy. Beyond final performance, we conduct extensive empirical analyses of RM-R1, including studies of its scaling effects, ablations of our training recipes, comparisons with non-reasoning baselines, training dynamics, and detailed case studies. In summary, our main contributions are as follows:
@@
  • We design a training recipe based on reasoning-oriented distillation and RL that produces a set of reward models – RM-R1 – that can outperform larger models by up to 4.9% on average.
@@
2 RM-R1
Figure 2: Training pipeline of RM-R1. Starting from an instruct model (GenRM), RM-R1 training involves two stages: Distillation and Reinforcement Learning (RL). In the Distillation stage, we use high-quality synthesized data to bootstrap RM-R1’s reasoning ability. In the RL stage, RM-R1’s reasoning ability for reward modeling is further strengthened. After distillation, a GenRM evolves into a ReasRM. RM-R1 further differentiates itself by being RL finetuned on preference data.
Figure 2 presents the overall training pipeline of RM-R1, which consists of two stages: reasoning distillation and reinforcement learning. (1) Reasoning Distillation: Starting from an off-the-shelf instruction-tuned model (e.g., Qwen-2.5-14B-Instruct), we further train the model using synthesized high-quality reasoning traces. This stage equips RM-R1 with essential reasoning capabilities required for effective reward modeling. (2) Reinforcement learning: While distillation is effective for injecting reasoning patterns, distilled models often overfit to specific patterns in the training data, limiting their generalization ability. To overcome this limitation, we introduce a reinforcement learning phase that further optimizes the model, resulting in the final version of RM-R1.
2.1 Task Definition
Given a preference dataset:
$\mathcal{D}=\{(x^{(i)},y_{a}^{(i)},y_{b}^{(i)},l^{(i)})\}_{i=1}^{N},$ (1)
where $x$ is a prompt, $y_{a}$ and $y_{b}$ are two different responses for $x$, and $l\in\{a,b\}$ is the ground truth label that indicates the preferred response. We define the generative reward modeling task as follows:
Let $r_{\theta}$ denote a generative reward model parameterized by $\theta$. For each data sample, $r_{\theta}$ generates a textual judgment $j$ consisting of ordered tokens $j=(j_{1},j_{2},\dots,j_{T})$, modeled by:
Note that $j$ contains $r_{\theta}$’s prediction of the preferred response $\hat{l}\subset j$. The overall objective is:
2.2 Reasoning Distillation for Reward Modeling
For an instruction-tuned model (e.g., Qwen-2.5-14b-instruct [35]), it is quite intuitive to turn it into a GenRM simply by prompting. However, without fine-tuning on reward modeling reasoning traces, these models may struggle to conduct consistent judgments. To bootstrap its reasoning potential, we start with training an instruction-tuned model with long reasoning traces synthesized for reward modeling. Specifically, we sample $M$ data samples from $\mathcal{D}$ and denote it as $\mathcal{D}_{\rm sub}$. Given a data sample $(x^{(i)},y_{a}^{(i)},y_{b}^{(i)},l^{(i)})\in\mathcal{D}_{\rm sub}$, we ask an “oracle” model like o3 or claude-3-7-sonnet to generate its structured reasoning trace $r^{(i)}$ justifying why $y_{l}^{(i)}$ is chosen as the correct answer of $x^{(i)}$. We then construct the reasoning trace ground truth:
$y_{\mathrm{trace}}^{(i)}=r^{(i)}\oplus l^{(i)},$ (3)
$\mathcal{D}_{\mathrm{distill}}=\{(x^{(i)},y_{\mathrm{trace}}^{(i)})\}_{i=1}^{M}.$ (4)
$\mathcal{L}_{\mathrm{distill}}(\theta)=-\sum_{(x,y)\in\mathcal{D}_{\mathrm{distill}}}\sum_{t=1}^{T}\log r_{\theta}\left(y_{t}\mid x,y_{<t}\right),$ (5)
where $y_{<t}=(y_{1},y_{2},...,y_{t-1})$ denotes the sequence of tokens preceding position $t$. More details of generating high-quality reasoning chains are included in Appendix B.
2.3 RL Training
Although distillation is a proper way to turn a general generative model into a GenRM, it often suffers from overfitting to certain patterns and constrains the model’s ability to generalize its reasoning abilities for critical thinking [9, 29] , which is essential for reward modeling. To address this, we propose to integrate RL as a more powerful learning paradigm to enhance the model’s ability to conduct reasoning-based rewarding. Training a policy model using RL has been widely seen in the preference optimization phase of LLM post-training [23], and it is quite natural to extend this paradigm for training a ReasRM. To be specific, we directly treat our reward model $r_{\theta}(j\mid x,y_{a},y_{b})$ as if it is a policy model:
where $r_{\mathrm{ref}}$ is the reference reward model. In practice, we use the checkpoint before training as $r_{\mathrm{ref}}$, and that means $r_{\mathrm{ref}}$ could be an off-the-shelf LLM or the LLM obtained after the distillation step in Section 2.2. $\mathcal{R}(x,j)$ is the reward function, and $\mathbb{D}_{\mathrm{KL}}$ is KL-divergence. The $x$ denotes input prompts drawn from the preference data $\mathcal{D}$. The $j$ indicates the text generated by the reward model, which includes the reasoning trace. In practice, we use Group Relative Policy Optimization (GRPO) [27] to optimize the objective in Equation 6, the details of which can be find in Appendix C.
2.3.1 Chain-of-Rubrics (CoR) Rollout
2. Assign weights to each rubric item based on their relative importance. 3. Inside <rubric>, include a <justify>…</justify> section explaining why you chose those rubric criteria and weights. 4. Compare both Chatbot responses according to the rubric. 5. Provide your evaluation inside <eval>…</eval> tags, using <quote_A>, <summary_A>, <quote_B>, and <summary_B> as described above. 6. End with your final judgment in the format: <answer>[[A]]</answer> or <answer>[[B]]</answer>
Figure 3: The system prompt used for RM-R1 rollout.
To facilitate the distilled models to proactively generate critical reasoning traces, we design a system prompt as shown in Figure 3 during rollout. Intuitively, reward modeling for general domain (e.g., chat, safety, etc.) and reasoning domain (e.g., math, code, etc.) should focus on different angles. For example, for the chat domain, we may care more about some aspects that can be expressed in textual rubrics (e.g., be polite), yet for the reasoning domain, we usually care more about logical coherence and answer correctness. Based on this intuition, we instruct ${r_{\theta}}$ to classify each preference data sample $\{(x,y_{c},y_{r})\}$ into one of the two <type>: Chat or Reasoning. For each <type>, we prompt ${r_{\theta}}$ to carry out the behavior corresponding to that type step by step: For reasoning tasks, we ask ${r_{\theta}}$ to solve $x$ on its own. During the <eval> phase, $r_{\theta}$ compares $y_{c}$ and $y_{r}$ conditioned on its own </solution> and selects an <answer>. Regarding the Chat type, we instead ask ${r_{\theta}}$ to think about and justify the <rubric> for grading the chat quality (including safety). Since Chat scenarios do not have a verifiable reward, we instruct the model to select the <answer> based on its own <rubric>.
2.3.2 Reward Design
Rule-based reward mechanisms have demonstrated strong empirical performance to facilitate reasoning [12]. In our training, we further simplify the reward formulation and merely focus on the correctness-based component, in line with prior works [27, 17].
Formally, our reward is defined as follows:
$\mathcal{R}(x,j|y_{a},y_{b})=\begin{cases}1&\text{if\ }\hat{l}=l,\\ -1&\text{otherwise.}\end{cases}$ (7)
where $\hat{l}$ is extracted from $j$, wrapped between the <answer> and </answer> tokens. We have also tried adding the format reward to the overall reward, but found that the task performance does not have a significant difference. The rationale behind only focusing on correctness is that the distilled models have already learned to follow instructions and format their responses properly.
3 Experiments
3.1 Experimental Setup
We evaluate RM-R1 on three primary benchmarks: RewardBench [16], RM-Bench [20], and RMB [39]. Our training set utilizes a cleaned subset of Skywork Reward Preference 80K [19], 8K examples from Code-Preference-Pairs, and the full Math-DPO-10K [15] dataset. For baselines, we compare RM-R1 with RMs from three main categories: ScalarRMs, GenRMs, and ReasRMs. Further details on the benchmarks, dataset construction, and specific baseline models are provided in Appendix D.
Table 1: The performance comparison between best-performing baselines. Bold numbers indicate the best performance, Underlined numbers indicate the second best. The DeepSeek-GRM models are not open-weighted, so we use the numbers on their tech report. The more detailed numbers on RewardBench, RM-Bench, and RMB are in Appendix Table 6,  Table 7, and Table 8
Models RewardBench RM-Bench RMB Average
ScalarRMs
SteerLM-RM-70B 88.8 52.5 58.2 66.5
Eurus-RM-7b 82.8 65.9 68.3 72.3
Internlm2-20b-reward 90.2 68.3 62.9 73.6
Skywork-Reward-Gemma-2-27B 93.8 67.3 60.2 73.8
Internlm2-7b-reward 87.6 67.1 67.1 73.9
ArmoRM-Llama3-8B-v0.1 90.4 67.7 64.6 74.2
Nemotron-4-340B-Reward 92.0 69.5 69.9 77.1
Skywork-Reward-Llama-3.1-8B 92.5 70.1 69.3 77.5
infly/INF-ORM-Llama3.1-70B 95.1 70.9 70.5 78.8
GenRMs
Claude-3-5-sonnet-20240620 84.2 61.0 70.6 71.9
Llama3.1-70B-Instruct 84.0 65.5 68.9 72.8
Gemini-1.5-pro 88.2 75.2 56.5 73.3
Skywork-Critic-Llama-3.1-70B 93.3 71.9 65.5 76.9
GPT-4o-0806 86.7 72.5 73.8 77.7
ReasRMs
JudgeLRM 75.2 64.7 53.1 64.3
DeepSeek-PairRM-27B 87.1 58.2
DeepSeek-GRM-27B-RFT 84.5 67.0
DeepSeek-GRM-27B 86.0 69.0
Self-taught-evaluator-llama3.1-70B 90.2 71.4 67.0 76.2
Our Methods
RM-R1-DeepSeek-Distilled-Qwen-7B 80.1 72.4 55.1 69.2
RM-R1-Qwen-Instruct-7B 85.2 70.2 66.4 73.9
RM-R1-Qwen-Instruct-14B 88.2 76.1 69.2 77.8
RM-R1-DeepSeek-Distilled-Qwen-14B 88.9 81.5 68.5 79.6
RM-R1-Qwen-Instruct-32B 91.4 79.1 73.0 81.2
RM-R1-DeepSeek-Distilled-Qwen-32B 90.9 83.9 69.8 81.5
3.2 Main Results
Table 1 compares the overall performance of RM-R1 with existing strongest baseline models. The more detailed numbers on RewardBench, RM-Bench, and RMB are in Table 6,  Table 7, and Table 8 in Appendix F. For the baselines, we reproduce the numbers if essential resources are open-sourced (e.g., model checkpoints, system prompts). Otherwise, we use the numbers reported in the corresponding tech report or benchmark leaderboard. For each benchmark, we select the best-performing models in each category for brevity. Our key findings are summarized below:
State-of-the-Art Performance. Our RM-R1-DeepSeek-Distilled-Qwen-14B model surpasses all previous leading Reward Models (RMs), including INF-ORM-Llama3.1-70B, Nemotron-4-340B-Reward, and GPT-4o, while operating at a much smaller scale on average. Our 32B models, RM-R1-Qwen-Instruct-32B and RM-R1-DeepSeek-Distilled-Qwen-32B, further extend this performance advantage by a non-trivial margin. The success of RM-R1 is attributable to both our meticulously designed training methodology and the effective scaling of our models, as extensively analyzed in Section 4.1 and Section 4.2. In particular, RM-R1 outperforms existing top-tier ScalarRMs. This highlights the considerable potential of ReasRMs, a category where prior GenRMs have exhibited suboptimal performance and are generally not comparable to their scalar counterparts. Simultaneously, our comprehensive evaluation indicates that the top-performing scalar models on RewardBench do not consistently achieve state-of-the-art (SOTA) performance; in fact, larger models frequently underperform smaller ones. This evaluation underscores the need for a more comprehensive and systematic approach to RM assessment.
Effective Training towards Reasoning for Reward Modeling. Our specialized, reasoning-oriented training pipeline delivers substantial performance gains. For instance, RM-R1-Qwen-Instruct-14B consistently surpasses Self-taught-evaluator-llama-3.1-70B, a reasoning model five times its size. The RM-R1 model series also demonstrates impressive results on RM-Bench, exceeding the top-performing baseline by up to 8.7%. On this most reasoning-intensive benchmark, RM-R1-DeepSeek-Distilled-Qwen-32B establishes a new state-of-the-art. It achieves 91.8% accuracy in math and 74.1% in code, outperforming the previous best models (73% in math and 63% in code) by significant margins. Furthermore, it also records the strongest reasoning performance among our released models on RewardBench. Despite its performance, our Instruct-based models are remarkably data-efficient, reaching competitive performance using only 8.7K examples for distillation—compared to the 800K examples used in training DeepSeek-Distilled [12]. Overall, our study underscores the significant potential of directly adapting large reasoning models into highly effective reward models.
4 Analysis
In this section, we present a series of empirical analyses to understand the key ingredients for training effective reasoning reward models. Our analysis spans scaling effects, design decisions, reasoning ablations, and a case study. We also present additional analysis on training dynamics in Section G.2.
4.1 Training Recipes
We first investigate the key ingredients underlying the successful training of RM-R1. Through a series of ablation studies, we examine our design choices to identify effective strategies for training high-quality reasoning reward models. We compare the following settings: Cold Start RL, Cold Start RL + Rubrics, Cold Start RL + Rubrics + Query Categorization (QC), and Distilled + RL + Rubrics + QC (i.e., RM-R1). The details of these settings are in Section G.1.
Table 2: Ablation study of the design choices for Reasoning Training on RewardBench.
Method Chat Chat Hard Safety Reasoning Average
Instruct (Original) 95.8 74.3 86.8 86.3 85.8
Instruct + Cold Start RL 92.5 81.5 89.7 94.4 89.5
Instruct + Cold Start RL + Rubrics 93.0 82.5 90.8 94.2 90.1
Instruct + Cold Start RL + Rubrics + QC 92.3 82.6 91.6 96.3 90.8
RM-R1 95.3 83.1 91.9 95.2 91.4
  • CoR prompting optimizes RM rollout and boosts reasoning performance. Instructing RM-R1 to self-generate chat rubrics or problem solutions before judgment helps overall performance, especially for chat and safety tasks. Incorporating explicit query categorization into the prompt notably improves reasoning performance, suggesting that clearer task guidance benefits learning.
[…]
4.2 Scaling Effects
We then investigate how model performance varies with scale, considering both model size and inference-time compute. In some cases – such as scalar-valued sequence models from Skywork [19] and InternLM2 [6] – the smaller models (8b/7b) outperforms the larger ones (27b/20b), showing no advantage of scaling. In this subsection, we show that this trend does not hold for RM-R1, where scaling brings clear and substantial improvements.
4.2.1 Model Sizes
We first analyze the impact of model scale. Our study is based on the Qwen-2.5-Instruct model family at three sizes: 7B, 14B, and 32B. We evaluate performance improvements resulting from our training procedure described in Section 2, with results averaged across three key benchmarks: RewardBench, RM-Bench, and RMB.
For each model size, we compare the original and post-training performance. Figure 4(a) plots the relative improvement (%) with respect to model size. Observing an approximately linear trend, we fit a linear regression model and extrapolate to hypothetical scales of 3B and 72B, shown using faint markers and dashed extensions. The results strongly support a scaling law for reasoning reward models: larger models not only result in an absolute better final performance but also consistently yield greater performance gains. This aligns with the intuition that our training effectively leverages the superior reasoning capabilities of larger models.
(a) Model Size
(b) Inference Compute
Figure 4: Scaling effect of RM-R1. (a) Larger models benefit more from reasoning training. (b) Longer reasoning chains improve RM performance.
Method RewardBench RM-Bench RMB Avg.
Full Data
Instruct + SFT 90.9 75.4 65.9 77.4
Instruct + Distilled + SFT 91.2 76.7 65.4 77.8
RM-R1 * 91.4 79.1 73.0 81.2
9k Data
Instruct + SFT 88.8 74.8 66.9 76.6
Instruct + Distilled* 89.0 76.3 72.0 79.2
Table 3: Comparison of reasoning-based training versus SFT across benchmarks. * indicates reasoning-based methods. Reasoning training consistently yields better performance.
4.2.2 Inference-time Computation
Next, we examine how model performance varies with different compute budgets measured in number of tokens allowed during inference. Since this is particularly relevant to reasoning-focused models, we fix our base model to DeepSeek-R1-Distill-Qwen-14B. We evaluate average performance across the three key benchmarks using a wide range of inference-time compute budgets: 512, 1024, 2048, 4096, and 8192 tokens.
To ensure a fair comparison, we match the training rollout budget to the inference budget in each setting (i.e., we allow a maximum of $x$ tokens during training for a compute budget of $x$ at inference). All models are trained using GRPO with identical datasets and hyperparameter configurations. Figure 4(b) shows the relationship between compute budget and performance. We observe a clear improvement trend as the inference budget increases. This highlights the benefits of long reasoning chains in reward modeling.
[…]
4.3 Effectiveness of Reasoning Training
We now analyze the impact of reasoning-based training. Here, we demonstrate that reasoning-based training can outperform answer-only approaches. We consider the following settings:
In summary, methods with “+ RM-R1” or “+ Distilled” represent reasoning-based approaches, while the remaining methods are purely non-reasoning-based approaches. In Table 3, we report the results measured across the three benchmarks. The findings clearly demonstrate that reasoning training significantly benefits reward model performance. Under fair comparisons (i.e., training on exactly the same amount of data), reasoning-based models consistently outperform their SFT-only counterparts. In particular, even high-quality distillation alone, applied to a small subset of the data, provides notable gains, highlighting the value of incorporating structured intermediate reasoning.
[…]
4.4 Case Study
Table 4: Case Study of RM-R1(succinct version). Complete model responses are shown in Table 9 and Table 10. Cold-Start RL evaluation tends to focus on superficial features (e.g., number of suggestions, relevance), whereas RM-R1 produces high-quality rubrics (e.g., accuracy for factual questions), faithfully follows them, and evaluates responses based on their actual content.
[…]
To gain deeper insights into RM-R1, we further conduct a case study using Qwen-14B-Instruct, comparing its behavior with a cold-start RL model. The results, summarized in Table 4, reveal the following key observations:
Transparent Judging Process.
High-Quality, Question-Dependent Rubrics.
Faithful Adherence to Rubrics and Content-Based Judgement.